# Concentration of discrepancy–based approximate Bayesian computation via Rademacher complexity

### v4

@@ Appendix B Proofs of Theorems, Corollaries and Propositions
Since $-n(\bar{\varepsilon}_{n}/3-2\mathfrak{R}_{n}(\mathfrak{F}))^{2}=-n\bar{\varepsilon}^{2}_{n}[1/9+4(\mathfrak{R}_{n}(\mathfrak{F})/\bar{\varepsilon}_{n})^{2}-(4/3)\mathfrak{R}_{n}(\mathfrak{F})/\bar{\varepsilon}_{n}]$, it follows that $-n(\bar{\varepsilon}_{n}/3-2\mathfrak{R}_{n}(\mathfrak{F}))^{2}\rightarrow-\infty$ when $n\to\infty$. From the above settings we also have that $n\bar{\varepsilon}^{2}_{n}\to~{}\infty$ and $\mathfrak{R}_{n}(\mathfrak{F})/\bar{\varepsilon}_{n}\to 0$, when $n\to\infty$. Therefore, as a consequence, we obtain $1-\exp[-n(\bar{\varepsilon}_{n}/3-2\mathfrak{R}_{n}(\mathfrak{F}))^{2}/2b^{2}]\rightarrow 1$ as $n\to\infty$. Hence, in the rest of this proof, we will restrict to the event $\{\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\mu^{*})\leq\bar{\varepsilon}_{n}/3\}$. Denote with $\mathbb{P}_{\theta,z_{1:n}}$ the joint distribution of $\theta\sim\pi$ and $z_{1:n}$ i.i.d. from $\mu_{\theta}$. By definition of conditional probability, for any $c_{2}$, including $c_{2}>2\mathfrak{R}_{n}(\mathfrak{F})$, it follows that
$\begin{split}&\pi_{n}^{(\varepsilon^{*}+\bar{\varepsilon}_{n})}\left(\{\theta:\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})>\varepsilon^{*}+4\bar{\varepsilon}_{n}/3+c_{2}\}\right)\\ &\qquad=\frac{\mathbb{P}_{\theta,z_{1:n}}\left[\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})>\varepsilon^{*}+4\bar{\varepsilon}_{n}/3+c_{2},\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})\leq\varepsilon^{*}+\bar{\varepsilon}_{n}\right]}{\mathbb{P}_{\theta,z_{1:n}}\left[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})\leq\varepsilon^{*}+\bar{\varepsilon}_{n}\right]}.\end{split}$ (B.4)
To derive an upper bound for the above ratio, we first identify an upper bound for its numerator. In addressing this goal, we leverage the triangle inequality $\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\mu_{\theta})+\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})+\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\mu^{*})$, since $\mathcal{D}_{\mathfrak{F}}$ is a semimetric, and the previously–proved result that the event $\{\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\mu^{*})\leq\bar{\varepsilon}_{n}/3\}$ has $\mathbb{P}_{y_{1:n}}$–probability going to $1$, thereby obtaining
@@
$\leq\mathbb{P}_{\theta,z_{1:n}}[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})+\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\mu_{\theta})+\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\mu^{*})>\varepsilon^{*}+4\bar{\varepsilon}_{n}/3+c_{2},$
$\qquad\qquad\ \mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})\leq\varepsilon^{*}+\bar{\varepsilon}_{n}]$
$\leq\mathbb{P}_{\theta,z_{1:n}}\left[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\mu_{\theta}){+}\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\mu^{*})>\bar{\varepsilon}_{n}/3{+}c_{2}\right]\leq\mathbb{P}_{\theta,z_{1:n}}\left[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\mu_{\theta})>c_{2}\right].$
Rewriting $\mathbb{P}_{\theta,z_{1:n}}\left[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\mu_{\theta})>c_{2}\right]$ as $\int_{\theta\in\Theta}\mathbb{P}_{z_{1:n}}\left[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\mu_{\theta})>c_{2}\mid\theta\right]\pi({\rm d}\theta)$ and applying (B.3) to $z_{1:n}$ yields
@@
Therefore, the numerator of the ratio in equation (B.4) can be upper bounded by $\exp[-n(c_{2}-2\mathfrak{R}_{n}(\mathfrak{F}))^{2}/2b^{2}]$ for any $c_{2}>2\mathfrak{R}_{n}(\mathfrak{F})$. As for the denominator, defining the event $\mbox{E}_{n}:=\{\theta\in\Theta:\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\varepsilon^{*}+\bar{\varepsilon}_{n}/3\}$ and applying again the triangle inequality, we have that
@@
where the last inequality follows directly from the fact that it is possible to restrict to the event $\{\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\mu^{*})\leq\bar{\varepsilon}_{n}/3\}$, and that we are integrating over $\mbox{E}_{n}:=\{\theta\in\Theta:\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\varepsilon^{*}+\bar{\varepsilon}_{n}/3\}$. Applying again Equation (B.3) to $z_{1:n}$, with $c_{1}=\bar{\varepsilon}_{n}/3>2\mathfrak{R}_{n}(\mathfrak{F})$, the last term of the above inequality can be further lower bounded by
@@
with $\pi(\mbox{E}_{n})\geq c_{\pi}(\bar{\varepsilon}_{n}/3)^{L}$ by (II), and, as shown before, $1-\exp[-n(\bar{\varepsilon}_{n}/3-2\mathfrak{R}_{n}(\mathfrak{F}))^{2}/2b^{2}]\to 1$, when $n\to\infty$, which implies, for $n$ large enough, $21-\exp[-n(\bar{\varepsilon}_{n}/3-2\mathfrak{R}_{n}(\mathfrak{F}))^{2}/2b^{2}]>1/$ 2. Leveraging both results, the denominator in (B.4) is lower bounded by $(c_{\pi}/2)\left(\bar{\varepsilon}_{n}/{3}\right)^{L}$. Let us now combine the upper and lower bounds derived, respectively, for the numerator and the denominator of the ratio in (B.4), to obtain
@@
with $\mathbb{P}_{y_{1:n}}$–probability going to $1$ as $n\rightarrow\infty$. To conclude the proof it suffices to replace $c_{2}$ in (B.5) with $2\mathfrak{R}_{n}(\mathfrak{F})+\sqrt{(2b^{2}/n)\log(M_{n}/\bar{\varepsilon}_{n}^{L})}$, which is never lower than $2\mathfrak{R}_{n}(\mathfrak{F})$. Finally, setting $M_{n}=n$ yields the statement of Theorem 3.2. ∎
@@
Corollary 3.2 follows by replacing the bounds in the proof of Corollary 1 by Bernton et al. (2019) with the newly–derived ones in Theorem 3.2. ∎
@@
Recall that in the case of mmd with bounded kernels we have $\mathfrak{R}_{n}(\mathfrak{F})\leq n^{-1/2}$. Hence, for the upper and lower bounds on $p_{n}$ in (3.1) it holds
Combining the above inequalities with the result in (3.1), and taking the limit for $n\rightarrow\infty$, proves the first part of the statement. The second part is a direct application of Corollary 3.1 to the case of mmd with bounded kernels, after noticing that the aforementioned inequality $\mathfrak{R}_{n}(\mathfrak{F})\leq n^{-1/2}$ implies $\mathfrak{R}_{n}(\mathfrak{F})\rightarrow 0$ as $n\rightarrow\infty$. ∎
Proof of Corollary 4.2.
To prove Corollary 4.2, it suffices to plug $\bar{\varepsilon}_{n}=[(\log n)/n]^{\frac{1}{2}}$ and ${b=1}$ into the statement of Theorem 3.2, and then upper–bound the resulting radius via the inequalities $\mathfrak{R}_{n}(\mathfrak{F})\leq n^{-1/2}$ and $\log n\geq 1$. The latter holds for any $n\geq 3$ and hence for $n\rightarrow\infty$. ∎
@@
We first show that, under $(A1)$$(A3)$, Assumptions 1–2 in Bernton et al. (2019) are satisfied when $f_{n}(\bar{\varepsilon}_{n})=1/(n\bar{\varepsilon}^{2}_{n})$ and $c(\theta)=\mathbb{E}_{z}\left[k(z,z)\right]$, with $z\sim\mu_{\theta}$. To this end, first recall that, by standard properties of mmd,
$\mathcal{D}^{2}_{\textsc{mmd}}(\mu_{1},\mu_{2})=\mathbb{E}_{x_{1},x_{1}^{\prime}}\left[k(x_{1},x_{1}^{\prime})\right]-2\mathbb{E}_{x_{1},x_{2}}\left[k(x_{1},x_{2})\right]+\mathbb{E}_{x_{2},x_{2}^{\prime}}\left[k(x_{2},x_{2}^{\prime})\right],$ (B.6)
with $x_{1},x_{1}^{\prime}\sim\mu_{1}$ and $x_{2},x_{2}^{\prime}\sim\mu_{2}$, all independently; see e.g., Chérief-Abdellatif and Alquier (2022). Since $k(x,x^{\prime})=\left<\phi(x),\phi(x^{\prime})\right>_{\mathcal{H}}$ (e.g., Muandet et al., 2017), the above result implies
$\begin{split}&\mathcal{D}^{2}_{\textsc{mmd}}(\mu_{1},\mu_{2})\\ &=\mathbb{E}_{x_{1},x_{1}^{\prime}}\left[\left<\phi(x_{1}),\phi(x_{1}^{\prime})\right>_{\mathcal{H}}\right]-2\mathbb{E}_{x_{1},x_{2}}\left[\left<\phi(x_{1}),\phi(x_{2})\right>_{\mathcal{H}}\right]+\mathbb{E}_{x_{2},x_{2}^{\prime}}\left[\left<\phi(x_{2}),\phi(x_{2}^{\prime})\right>_{\mathcal{H}}\right]\\ &=||\mathbb{E}_{x_{1}}[\phi(x_{1})]||_{\mathcal{H}}^{2}-2\left<\mathbb{E}_{x_{1}}[\phi(x_{1})],\mathbb{E}_{x_{2}}[\phi(x_{2})]\right>_{\mathcal{H}}+||\mathbb{E}_{x_{2}}[\phi(x_{2})]||_{\mathcal{H}}^{2}\\ &=\left\|\mathbb{E}_{x_{1}}\left[\phi(x_{1})\right]-\mathbb{E}_{x_{2}}\left[\phi(x_{2})\right]\right\|^{2}_{\mathcal{H}}.\end{split}$ (B.7)
Leveraging Equations (B.6)–(B.7) and basic Markov inequalities, for any $\bar{\varepsilon}_{n}\geq 0$, it holds
@@
with $f_{n}(u)=1/(nu^{2})$ strictly decreasing in $u$ for any fixed $n$, and $f_{n}(u)\rightarrow 0$ as $n\rightarrow\infty$, for fixed $u$. Moreover, by Assumptions $(A2)$$(A3)$, $c(\theta)=\mathbb{E}_{z}\left[k(z,z)\right]$ is $\pi$–integrable and there exist a $\delta_{0}>0$ and a $c_{0}>0$ such that $c(\theta)<c_{0}$ for any $\theta$ satisfying $(\mathbb{E}_{z,z^{\prime}}\left[k(z,z^{\prime})\right]-2\mathbb{E}_{z,y}\left[k(y,z)\right]+\mathbb{E}_{y,y^{\prime}}\left[k(y,y^{\prime})\right])^{1/2}=\mathcal{D}_{\textsc{mmd}}(\mu_{\theta},\mu^{*})\leq\varepsilon^{*}+\delta_{0}$. This ensures that Assumption 2 in Bernton et al. (2019) holds.
Under the assumptions in Proposition 4.1 it is, therefore, possible to apply Proposition 3 in Bernton et al. (2019) with $f_{n}(\bar{\varepsilon}_{n})=1/(n\bar{\varepsilon}^{2}_{n})$, $c(\theta)=\mathbb{E}_{z}\left[k(z,z)\right]$ and $R=M_{n}$, which yields the concentration result in Proposition 4.1. ∎
Proof of Lemma 6.1.
Let $(x_{t})_{t\in\mathbb{Z}}$ be a stochastic process with $\beta$–mixing coefficients $\beta(k)$, for $k\in\mathbb{N}$. Moreover, denote with $\mu^{(n)}$ and $\mu$ the joint distribution of a sample $x_{1:n}$ from the process $(x_{t})_{t\in\mathbb{Z}}$ and its constant marginal $\mu=\mu^{(1)}$, respectively. Then, by combining Proposition 2 and Lemma 2 in Mohri & Rostamizadeh (2008) under our notation, we have that for any $b$–uniformly bounded class $\mathfrak{F}$, any integer $n\geq 1$ and any scalar $\delta\geq 0$, the inequality

### v5

@@ Appendix D Proofs of Theorems, Corollaries and Propositions
Since $-n(\bar{\varepsilon}_{n}/3-2\mathfrak{R}_{n}(\mathfrak{F}))^{2}=-n\bar{\varepsilon}^{2}_{n}[1/9+4(\mathfrak{R}_{n}(\mathfrak{F})/\bar{\varepsilon}_{n})^{2}-(4/3)\mathfrak{R}_{n}(\mathfrak{F})/\bar{\varepsilon}_{n}]$, it follows that $-n(\bar{\varepsilon}_{n}/3-2\mathfrak{R}_{n}(\mathfrak{F}))^{2}\rightarrow-\infty$ when $n\to\infty$. From the above settings we also have that $n\bar{\varepsilon}^{2}_{n}\to~{}\infty$ and $\mathfrak{R}_{n}(\mathfrak{F})/\bar{\varepsilon}_{n}\to 0$, when $n\to\infty$. Therefore, as a consequence, we obtain $1-\exp[-n(\bar{\varepsilon}_{n}/3-2\mathfrak{R}_{n}(\mathfrak{F}))^{2}/2b^{2}]\rightarrow 1$ as $n\to\infty$. Hence, in the rest of this proof, we will restrict to the event $\{\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\mu^{*})\leq\bar{\varepsilon}_{n}/3\}$.
Denote with $\mathbb{P}_{\theta,z_{1:n}}$ the joint distribution of $\theta\sim\pi$ and $z_{1:n}$ i.i.d. from $\mu_{\theta}$. By definition of conditional probability, for any $c_{2}$, including $c_{2}>2\mathfrak{R}_{n}(\mathfrak{F})$, it follows that
$\begin{split}&\pi_{n}^{(\varepsilon^{*}+\bar{\varepsilon}_{n})}\left(\{\theta:\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})>\varepsilon^{*}+4\bar{\varepsilon}_{n}/3+c_{2}\}\right)\\ &\qquad=\frac{\mathbb{P}_{\theta,z_{1:n}}\left[\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})>\varepsilon^{*}+4\bar{\varepsilon}_{n}/3+c_{2},\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})\leq\varepsilon^{*}+\bar{\varepsilon}_{n}\right]}{\mathbb{P}_{\theta,z_{1:n}}\left[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})\leq\varepsilon^{*}+\bar{\varepsilon}_{n}\right]}.\end{split}$ (D.4)
To derive an upper bound for the above ratio, we first identify an upper bound for its numerator. In addressing this goal, we leverage the triangle inequality $\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\mu_{\theta})+\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})+\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\mu^{*})$, since $\mathcal{D}_{\mathfrak{F}}$ is a semimetric, and the previously-proved result that the event $\{\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\mu^{*})\leq\bar{\varepsilon}_{n}/3\}$ has $\mathbb{P}_{y_{1:n}}$–probability going to $1$, thereby obtaining
@@
$\qquad\quad\leq\mathbb{P}_{\theta,z_{1:n}}[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})+\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\mu_{\theta})+\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\mu^{*})>\varepsilon^{*}+4\bar{\varepsilon}_{n}/3+c_{2},$
$\qquad\quad\qquad\qquad\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})\leq\varepsilon^{*}+\bar{\varepsilon}_{n}]$
$\qquad\quad\leq\mathbb{P}_{\theta,z_{1:n}}\left[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\mu_{\theta})+\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\mu^{*})>\bar{\varepsilon}_{n}/3+c_{2}\right]\leq\mathbb{P}_{\theta,z_{1:n}}\left[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\mu_{\theta})>c_{2}\right].$
Rewriting $\mathbb{P}_{\theta,z_{1:n}}\left[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\mu_{\theta})>c_{2}\right]$ as $\int_{\theta\in\Theta}\mathbb{P}_{z_{1:n}}\left[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\mu_{\theta})>c_{2}\mid\theta\right]\pi({\rm d}\theta)$ and applying (D.3) to $z_{1:n}$ yields
@@
Hence, the numerator of the ratio in Equation (D.4) can be upper bounded by $\exp[-n(c_{2}-2\mathfrak{R}_{n}(\mathfrak{F}))^{2}/2b^{2}]$ for any $c_{2}>2\mathfrak{R}_{n}(\mathfrak{F})$. As for the denominator, defining the event $\mbox{E}_{n}:=\{\theta\in\Theta:\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\varepsilon^{*}+\bar{\varepsilon}_{n}/3\}$ and applying again the triangle inequality, we have that
@@
where the last inequality follows directly from the fact that it is possible to restrict to the event $\{\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\mu^{*})\leq\bar{\varepsilon}_{n}/3\}$, and that we are integrating over $\mbox{E}_{n}:=\{\theta\in\Theta:\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\varepsilon^{*}+\bar{\varepsilon}_{n}/3\}$. Applying again (D.3) to $z_{1:n}$, with $c_{1}=\bar{\varepsilon}_{n}/3>2\mathfrak{R}_{n}(\mathfrak{F})$, the last term of the above inequality can be further lower bounded by
@@
with $\pi(\mbox{E}_{n})\geq c_{\pi}(\bar{\varepsilon}_{n}/3)^{L}$ by (II), and, as shown before, $1-\exp[-n(\bar{\varepsilon}_{n}/3-2\mathfrak{R}_{n}(\mathfrak{F}))^{2}/2b^{2}]\to 1$, when $n\to\infty$, which implies, for $n$ large enough, $21-\exp[-n(\bar{\varepsilon}_{n}/3-2\mathfrak{R}_{n}(\mathfrak{F}))^{2}/2b^{2}]>1/$ 2. Leveraging both results, the denominator in (D.4) is lower bounded by $(c_{\pi}/2)\left(\bar{\varepsilon}_{n}/{3}\right)^{L}$. Let us now combine the upper and lower bounds derived, respectively, for the numerator and the denominator of the ratio in (D.4), to obtain
@@
with $\mathbb{P}_{y_{1:n}}$–probability going to $1$ as $n\rightarrow\infty$. To conclude the proof it suffices to replace $c_{2}$ in (D.5) with $2\mathfrak{R}_{n}(\mathfrak{F})+\sqrt{(2b^{2}/n)\log(M_{n}/\bar{\varepsilon}_{n}^{L})}$, which is never lower than $2\mathfrak{R}_{n}(\mathfrak{F})$. Finally, setting $M_{n}=n$ yields the statement of Theorem 3.3. ∎
Proof of Corollary B.1.
Corollary B.1 follows by replacing the bounds in the proof of Corollary 1 by Bernton et al. (2019) with the newly-derived ones in Theorem 3.3. ∎
@@
Recall that in the case of mmd with kernels bounded by $1$ we have $\mathfrak{R}_{n}(\mathfrak{F})\leq n^{-1/2}$. Hence, regarding the upper and lower bounds on $p_{n}$ in (3) it holds
Combining the above inequalities with the result in (3), and taking the limit for $n\rightarrow\infty$, proves the first part of the statement. The second part is a direct application of Corollary 3.2 to the case of mmd with bounded kernels, after noticing that the aforementioned inequality $\mathfrak{R}_{n}(\mathfrak{F})\leq n^{-1/2}$ implies $\mathfrak{R}_{n}(\mathfrak{F})\rightarrow 0$ as $n\rightarrow\infty$. ∎
@@
To prove Corollary 4.2, it suffices to plug $\bar{\varepsilon}_{n}=[(\log n)/n]^{\frac{1}{2}}$ and ${b=1}$ into the statement of Theorem 3.3, and then upper-bound the resulting radius via the inequalities $\mathfrak{R}_{n}(\mathfrak{F})\leq n^{-1/2}$ and $\log n\geq 1$. The latter holds for any $n\geq 3$ and hence for $n\rightarrow\infty$. ∎
@@
We first show that, under $(A1)$$(A3)$, Assumptions 1 and 2 made in Bernton et al. (2019) are satisfied under mmd when $f_{n}(\bar{\varepsilon}_{n})=1/(n\bar{\varepsilon}^{2}_{n})$ and $c(\theta)=\mathbb{E}_{z}\left[k(z,z)\right]$, with $z\sim\mu_{\theta}$.
Consistent with the above goal, first recall that, by standard properties of mmd,
$\mathcal{D}^{2}_{\textsc{mmd}}(\mu_{1},\mu_{2})=\mathbb{E}_{x_{1},x_{1}^{\prime}}\left[k(x_{1},x_{1}^{\prime})\right]-2\mathbb{E}_{x_{1},x_{2}}\left[k(x_{1},x_{2})\right]+\mathbb{E}_{x_{2},x_{2}^{\prime}}\left[k(x_{2},x_{2}^{\prime})\right],$ (D.6)
with $x_{1},x_{1}^{\prime}\sim\mu_{1}$ and $x_{2},x_{2}^{\prime}\sim\mu_{2}$, all independently; see e.g., Chérief-Abdellatif and Alquier (2022). Since $k(x,x^{\prime})=\left<\phi(x),\phi(x^{\prime})\right>_{\mathcal{H}}$ (see e.g., Muandet et al., 2017), the above result implies that
$\begin{split}&\mathcal{D}^{2}_{\textsc{mmd}}(\mu_{1},\mu_{2})\\ &=\mathbb{E}_{x_{1},x_{1}^{\prime}}\left[\left<\phi(x_{1}),\phi(x_{1}^{\prime})\right>_{\mathcal{H}}\right]-2\mathbb{E}_{x_{1},x_{2}}\left[\left<\phi(x_{1}),\phi(x_{2})\right>_{\mathcal{H}}\right]+\mathbb{E}_{x_{2},x_{2}^{\prime}}\left[\left<\phi(x_{2}),\phi(x_{2}^{\prime})\right>_{\mathcal{H}}\right]\\ &=||\mathbb{E}_{x_{1}}[\phi(x_{1})]||_{\mathcal{H}}^{2}-2\left<\mathbb{E}_{x_{1}}[\phi(x_{1})],\mathbb{E}_{x_{2}}[\phi(x_{2})]\right>_{\mathcal{H}}+||\mathbb{E}_{x_{2}}[\phi(x_{2})]||_{\mathcal{H}}^{2}\\ &=\left\|\mathbb{E}_{x_{1}}\left[\phi(x_{1})\right]-\mathbb{E}_{x_{2}}\left[\phi(x_{2})\right]\right\|^{2}_{\mathcal{H}}.\end{split}$ (D.7)
Leveraging Equations (D.6)–(D.7) and basic Markov inequalities, for any $\bar{\varepsilon}_{n}\geq 0$, it holds
@@
with $f_{n}(u)=1/(nu^{2})$ strictly decreasing in $u$ for any fixed $n$, and $f_{n}(u)\rightarrow 0$ as $n\rightarrow\infty$, for fixed $u$. Moreover, by Assumptions $(A2)$$(A3)$, $c(\theta)=\mathbb{E}_{z}\left[k(z,z)\right]$ is $\pi$-integrable and there exist a $\delta_{0}>0$ and a $c_{0}>0$ such that $c(\theta)<c_{0}$ for any $\theta$ satisfying $(\mathbb{E}_{z,z^{\prime}}\left[k(z,z^{\prime})\right]-2\mathbb{E}_{z,y}\left[k(y,z)\right]+\mathbb{E}_{y,y^{\prime}}\left[k(y,y^{\prime})\right])^{1/2}=\mathcal{D}_{\textsc{mmd}}(\mu_{\theta},\mu^{*})\leq\varepsilon^{*}+\delta_{0}$. This ensures that Assumption 2 in Bernton et al. (2019) holds.
Finally, note that Assumption 3 in Bernton et al. (2019), is verified by our Assumption (II). Therefore, under the assumptions in Proposition 4.3 it is possible to apply Proposition 3 in Bernton et al. (2019) with $f_{n}(\bar{\varepsilon}_{n})=1/(n\bar{\varepsilon}^{2}_{n})$, $c(\theta)=\mathbb{E}_{z}\left[k(z,z)\right]$ and $R=M_{n}$, which yields the concentration result in Proposition 4.3. ∎
Proof of Proposition 4.4.
Assumptions (A1’) and (A2’) imply that the norm $\|\|x\|\|_{\psi_{1}}$ (see Equation (14) in Lei (2020) for a definition) is uniformly bounded when $x\sim\mu_{\theta}$ and $x\sim\mu^{*}$. Thus, we can use (15) in Lei (2020) to obtain
$\mathbb{P}_{z_{1:n}}\left(\mathcal{D}_{\textrm{wass}}(\mu_{\theta},\hat{\mu}_{z_{1:n}})>u\right)\leq\exp[-c^{\prime}n(u-c_{1}n^{-1/\max(d,3)})_{+}^{2}]=:f_{n}(u)$
and, similarly, $\mathbb{P}_{y_{1:n}}\left(\mathcal{D}_{\textrm{wass}}(\mu^{*},\hat{\mu}_{y_{1:n}})>u\right)\leq f_{n}(u)$, where $c^{\prime}$ depends only on the constant $c$ in (A1’) and (A2’), and $c_{1}$ is a universal constant. Notice that (15) in Lei (2020) requires $d>2$. If $d=1$, we can define $x^{\prime}=(x,0,0)\in\mathbb{R}^{3}$ and apply the result in $\mathbb{R}^{3}$ (we can proceed similarly if $d=2$). This is why Proposition 4.4 is stated with $\max(d,3)$. The above bounds ensure that Assumptions 1–2 in Bernton et al. (2019) are met. Moreover, condition (II) verifies Assumption 3 in Bernton et al. (2019). Thus, we can apply Proposition 3 of Bernton et al. (2019) with $f_{n}(\cdot)$ defined as above and under the vanishing conditions on $\bar{\varepsilon}_{n}$ in Proposition 4.4. This implies that when $n\rightarrow\infty$, and $\bar{\varepsilon}_{n}\rightarrow 0$ such that $f_{n}(\bar{\varepsilon}_{n})\rightarrow 0$, for some ${C\in(0,\infty)}$ and any $M_{n}\in(0,\infty)$, with $\mathbb{P}_{y_{1:n}}$–probability going to $1$ as $n\rightarrow\infty$, it holds
$\pi_{n}^{(\varepsilon^{*}+\bar{\varepsilon}_{n})}(\{\theta:\mathcal{D}_{\textrm{wass}}(\mu_{\theta},\mu^{*})>\varepsilon^{*}+4\bar{\varepsilon}_{n}/3+f_{n}^{-1}\left(\bar{\varepsilon}_{n}^{L}/M_{n}\right)\})\leq C/M_{n}.$
Recall that our restriction $n^{-1/\max(d,3)}\ll\bar{\varepsilon}_{n}$, together with $n\bar{\varepsilon}^{2}_{n}\to\infty$, imply that for $n$ large enough $f_{n}(\cdot)$ is invertible and $f_{n}(\bar{\varepsilon}_{n})\rightarrow 0$. On such a range, we have that $f_{n}^{-1}\left(\bar{\varepsilon}_{n}^{L}/M_{n}\right)=[(1/(c^{\prime}n))\log(M_{n}/\bar{\varepsilon}_{n}^{L})]^{1/2}+c_{1}n^{-{1}/{\max(d,3)}},$ which concludes the proof. ∎
Proof of Lemma C.2.
Let $(x_{t})_{t\in\mathbb{Z}}$ be a stochastic process with $\beta$-mixing coefficients $\beta(k)$, for $k\in\mathbb{N}$. Moreover, denote with $\mu^{(n)}$ and $\mu$ the joint distribution of a sample $x_{1:n}$ from $(x_{t})_{t\in\mathbb{Z}}$ and its constant marginal $\mu=\mu^{(1)}$, respectively. Then, by combining Proposition 2 and Lemma 2 in Mohri and Rostamizadeh (2008) under our notation, we have that for any $b$-uniformly bounded class $\mathfrak{F}$, any integer $n\geq 1$ and any scalar $\delta\geq 0$, the inequality